# Searching for light long-lived particles at SHiNESS

### v1

@@ Abstract
Recently Soleti et al. [JHEP03(2024)148] proposed a new experiment called SHiNESS at the upcoming European Spallation Source (ESS) facility, making use of the 2-GeV proton beam there impinging on a fixed target, in order to search for hidden sterile neutrinos that could lie in different mass ranges and arise with distinct signatures. Such signatures include excesses in electron-positron pairs that may originate from displaced decays of long-lived particles (LLPs). At the ESS, the dominant sources of such LLPs are decays of $\pi^{+}$ mesons at rest and decays of $\mu^{+}$ leptons. We choose to investigate theoretical scenarios of long-lived light binos in the R-parity-violating supersymmetry and long-lived weak-violating electrophilic axionlike particles, as these LLPs can be produced from decays of $\pi^{+}$ and $\mu^{+}$. Since the $\pi^{+}$’s decay at rest at the ESS, we compute the spectra of the therefrom produced $\mu^{+}$’s or LLPs, and thus estimate the expected sensitivity reach of SHiNESS to the LLPs in these two scenarios. Our calculation shows that in most of the relevant benchmark scenarios, SHiNESS can probe large parameter regions of these models beyond the existing bounds, in just a couple of years of data-collection time.
@@ 3 Experimental setup
At the upcoming ESS facility, a 2-GeV proton beam hits a rotating tungsten target, producing large numbers of neutrons, serving the facility’s main purpose Garoby:2017vew . In total, $2.8\times 10^{23}$ protons on target are planned for each calendar year. In such proton-nucleus collisions, an enormous number of charged pions among other hadrons are produced. While negatively charged pions are absorbed by the nucleus before decaying, positively charged ones lose energy while they are traveling and finally decay at rest, leading to the following decay chain:
@@
Simulation indicates a neutrino yield of 0.3 per flavor ($\nu_{\mu},\nu_{e},$ and $\bar{\nu}_{\mu}$) per proton, implying $8.5\times 10^{22}$ neutrinos per flavor per year produced from $\pi^{+}$ decays at rest Baxter:2019mcx . Since the $\pi^{+}$ mesons decay to $\mu^{+}\nu_{\mu}$ with a branching fraction of $\sim 99.99\%$ ParticleDataGroup:2022pth , this corresponds to $8.5\times 10^{22}$ $\pi^{+}$ mesons produced each year and decaying at rest.
@@
We note that despite the target wheel being 2.6-m wide Garoby:2017vew and the non-prompt nature of both $\pi^{+}$ and $\mu^{+}$, we have assumed above, for simplicity, that the target is point-like, and that both the $\pi^{+}$ mesons and $\mu^{+}$ leptons decay at rest at the target. We expect this simplification of the setup to affect the final results to a rather minor extent (see discussion in Ref. Baxter:2019mcx ). Moreover, for the LLPs produced from $\mu^{+}$ decays, we assume that the kinematics of the $\mu^{+}$ leptons are completely determined by the two-body decays of the $\pi^{+}$ mesons at rest, with no interference from the experimental facility.
@@ 4.1 Light binos
For benchmarks B1.1 and B1.2, strong bounds exist on the single coupling $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the light bino’s mass, from $0\nu\beta\beta$ decays and HNL searches, cf. the discussion in Sec. 2. Additionally, for B1.1, the product bound given in Eq. (8) constrains the scenario tightly. Therefore, we choose to present only a sensitivity plot in Fig. 2 for B1.2, shown in the plane $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$, where we fix the light bino’s mass at 20 MeV (black), 70 MeV (scarlet), and 120 MeV (green). The horizontal dashed curves are current upper bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$, where we have picked, for each bino mass choice, the stronger limit between that derived from $0\nu\beta\beta$ decays and that from HNL searches. The vertical red curves are present upper limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ for $m_{\tilde{f}}=1$ TeV and 2 TeV (see Eq. (5)) which do not depend on $m_{\tilde{\chi}^{0}_{1}}$. We find that for all three choices of the bino mass, a triangle-shaped window that is still allowed in the shown parameter space can be excluded by SHiNESS in two years.
@@
Figure 3: Sensitivity reach of SHiNESS to the benchmark B3.1, displayed in the plane $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$. The red dashed horizontal curves are current limits on $\lambda_{121}/m^{2}_{\tilde{f}}$, cf. Eq. (4), for $m_{\tilde{f}}=1$ TeV and 2 TeV, and the gray dashed curve is the bound from recasting HNL searches, extracted from Ref. Dreiner:2023gir .
@@
For B4.1, the product bound as given in Eq. (7) is so strong that it has excluded all the parameter space that would be probed by SHiNESS if we set $\lambda_{121}=\lambda_{122}$. If the two non-vanishing $\lambda$ couplings are independent, when $\lambda_{121}/m^{2}_{\tilde{f}}$ is sufficiently small, SHiNESS can only probe $\lambda_{122}/m^{2}_{\tilde{f}}$ values that are already excluded by the product bound. For larger values of $\lambda_{121}/m^{2}_{\tilde{f}}$, the search is insensitive to $\lambda_{122}$ as the single coupling $\lambda_{121}$ leading to both production and decay of the light bino alone would result in sufficiently many signal events (see Fig. 3).
@@ 5 Conclusions
The focus of this work is on LLPs produced in rare decays at rest of $\pi^{+}$ mesons that arise from the 2-GeV proton beam at the facility hitting a rotating tungsten target and lose energy as they propagate. Specifically, we choose two theoretical models predicting light, MeV-scale LLPs, namely, a light bino in the RPV-SUSY, and a weak-violating electrophilic ALP. We consider production modes of the light bino including decays of $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ and $\mu^{+}\to e^{+}\overset{(-)}{\nu}\tilde{\chi}^{0}_{1}$ where $\mu^{+}$ stems from the $\pi^{+}\to\mu^{+}\nu_{\mu}$ decay at rest. The light bino then travels and decays inside SHiNESS into an electron-positron pair in association with a neutrino. Various benchmark scenarios, each with two non-vanishing RPV couplings mediating the production and decay of the light bino, respectively, are studied. Numerical results are then obtained and presented in the plane of either the production coupling vs. the decay coupling, or the RPV couplings (assumed to be equal) versus the light bino’s mass. While for some benchmark scenarios we find SHiNESS is unable to probe unexcluded parameter space of the model for reasons of kinematics or too strong current bounds, for other benchmarks SHiNESS is predicted to be able to exclude large, new territories in the parameter space with just two years’ data collection.
@@ Acknowledgements
We thank Martin Schürmann for useful discussions. W.L. is supported by National Natural Science Foundation of China (Grant No. 12205153). Y.Z. is supported by the National Natural Science Foundation of China under Grant No. 11805001 and the Fundamental Research Funds for the Central Universities under Grant No. JZ2023HGTB0222

### v2

@@ Abstract
Recently Soleti et al. [JHEP03(2024)148] proposed a new experiment called SHiNESS at the upcoming European Spallation Source (ESS) facility, making use of the 2-GeV proton beam there impinging on a fixed target, in order to search for hidden sterile neutrinos that could lie in different mass ranges and arise with distinct signatures. Such signatures include excesses in electron-positron pairs that may originate from displaced decays of long-lived particles (LLPs). At the ESS, the dominant sources of such LLPs are decays at rest of $\pi^{+}$ mesons and $\mu^{+}$ leptons. We choose to investigate theoretical scenarios of long-lived light binos in the R-parity-violating supersymmetry and long-lived weak-violating electrophilic axionlike particles, as these LLPs can be produced from decays of $\pi^{+}$ and $\mu^{+}$. Since the $\pi^{+}$’s and $\mu^{+}$’s decay at rest at the ESS, we compute the spectra of the therefrom produced LLPs, and thus estimate the expected sensitivity reach of SHiNESS to the LLPs in these two scenarios. Our calculation shows that in most of the relevant benchmark scenarios, SHiNESS can probe large parameter regions of these models beyond the existing bounds, in just a couple of years of data-collection time.
@@ 3 Experimental setup
At the upcoming ESS facility, a 2-GeV proton beam hits a rotating tungsten target, producing large numbers of neutrons, serving the facility’s main purpose Garoby:2017vew . In total, $2.8\times 10^{23}$ protons on target are planned for each calendar year. In such proton-nucleus collisions, an enormous number of charged pions among other hadrons are produced. While negatively charged pions are absorbed by the nucleus before decaying, positively charged ones lose energy via ionization effects while they are traveling and finally decay at rest, leading to the following decay chain:
@@
Simulation indicates a neutrino yield of 0.3 per flavor ($\nu_{\mu},\nu_{e},$ and $\bar{\nu}_{\mu}$) per proton, implying $8.5\times 10^{22}$ neutrinos per flavor per year produced from $\pi^{+}$ decays at rest Baxter:2019mcx . Since the $\pi^{+}$ mesons decay to $\mu^{+}\nu_{\mu}$ with a branching fraction of $\sim 99.99\%$ ParticleDataGroup:2022pth , this corresponds to $8.5\times 10^{22}$ $\pi^{+}$ mesons produced each year and decaying at rest. The $\mu^{+}$ leptons produced from $\pi^{+}$ decays also undergo ionization processes and quickly decay at rest in the close vicinity Baxter:2019mcx ; Burman:2003ek .
@@
We note that despite the target wheel being 2.6-m wide Garoby:2017vew , we have assumed above, for simplicity, that the target is point-like. Moreover, both $\pi^{+}$ mesons and $\mu^{+}$ leptons are assumed to decay at rest at the interaction point on the target, considering the ionization effects as mentioned above. We expect this simplification of the setup to affect the final results to a rather minor extent (see discussion in Ref. Baxter:2019mcx ).
@@ 4.1 Light binos
For benchmarks B1.1 and B1.2, strong bounds exist on the single coupling $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the light bino’s mass, from $0\nu\beta\beta$ decays and HNL searches, cf. the discussion in Sec. 2. Additionally, for B1.1, the product bound given in Eq. (8) constrains the scenario tightly. Therefore, we choose to present only a sensitivity plot in Fig. 2 for B1.2, shown in the plane $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$, where we fix the light bino’s mass at 20 MeV (black), 70 MeV (scarlet), and 120 MeV (green). The horizontal dashed curves are current upper bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$, where we have picked, for each bino-mass choice, the stronger limit between that derived from $0\nu\beta\beta$ decays and that from HNL searches. The vertical red curves are present upper limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ for $m_{\tilde{f}}=1$ TeV and 2 TeV (see Eq. (5)) which do not depend on $m_{\tilde{\chi}^{0}_{1}}$. We find that for all three choices of the bino mass, a triangle-shaped window that is still allowed in the shown parameter space can be excluded by SHiNESS in two years.
@@
Figure 3: Sensitivity reach of SHiNESS to the benchmark B3.1, displayed in the plane $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$. The red dashed horizontal curves are current limits on $\lambda_{121}/m^{2}_{\tilde{f}}$, cf. Eq. (4), for $m_{\tilde{f}}=1$ TeV and 2 TeV, and the gray dashed curve is the bound extracted from Ref. Dreiner:2023gir .
@@
For B4.1, the product bound as given in Eq. (7) is so strong that it has excluded all the parameter space that would be probed by SHiNESS if we set $\lambda_{121}=\lambda_{122}$. If the two non-vanishing $\lambda$ couplings are independent, when $\lambda_{121}/m^{2}_{\tilde{f}}$ is sufficiently small, SHiNESS can only probe $\lambda_{122}/m^{2}_{\tilde{f}}$ values that are already excluded by the product bound. For larger values of $\lambda_{121}/m^{2}_{\tilde{f}}$, the search is insensitive to $\lambda_{122}$ as the single coupling $\lambda_{121}$ leading to both production and decay of the light bino alone would result in sufficiently many signal events (see Fig. 3). For these reasons, we do not present numerical results for B4.1.
@@ 5 Conclusions
The focus of this work is on LLPs produced in rare decays at rest of $\pi^{+}$ mesons that arise from the 2-GeV proton beam at the facility hitting a rotating tungsten target and lose energy as they propagate. Specifically, we choose two theoretical models predicting light, MeV-scale LLPs, namely, a light bino in the RPV-SUSY, and a weak-violating electrophilic ALP. We consider production modes of the light bino including decays of $\pi^{+}\to e^{+}\tilde{\chi}^{0}_{1}$ and $\mu^{+}\to e^{+}\overset{(-)}{\nu}\tilde{\chi}^{0}_{1}$, where $\mu^{+}$ stems from the $\pi^{+}\to\mu^{+}\nu_{\mu}$ decay at rest and also decays at rest immediately. The light bino then travels and decays inside SHiNESS into an electron-positron pair in association with a neutrino. Various benchmark scenarios, each with two non-vanishing RPV couplings mediating the production and decay of the light bino, respectively, are studied. Numerical results are then obtained and presented in the plane of either the production coupling vs. the decay coupling, or the RPV couplings (assumed to be equal) versus the light bino’s mass. While for some benchmark scenarios we find SHiNESS is unable to probe unexcluded parameter space of the model for reasons of kinematics or too strong current bounds, for other benchmarks SHiNESS is predicted to be able to exclude large, new territories in the parameter space with just two years’ data collection.
@@ Acknowledgements
We thank Liangwen Chen and Martin Schürmann for useful discussions. Y.Z. is supported by the National Natural Science Foundation of China under Grant No. 11805001 and the Fundamental Research Funds for the Central Universities under Grant No. JZ2023HGTB0222. W.L. is supported by National Natural Science Foundation of China (Grant No. 12205153).